# Uncovering flow and deformation regimes in the coupled fluid-solid vestibular system

### v1

@@ 3.1 Expanded solution
$w_{0}(r,s,t)=-\frac{1}{4}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)(a(s)^{2}-r^{2}),$ (20a)
$w_{1}(r,\theta,s,t)=-\frac{1}{4}\kappa\frac{\partial\bar{p}_{1}}{\partial s}(a(s)^{2}-r^{2})+\frac{1}{16}\left(f(t)-3\kappa\frac{\partial p_{0}}{\partial s}\right)r(r^{2}-a(s)^{2})\cos\theta.$ (20b)
Figure 4: Velocity profiles predicted by (20) as the torus aspect ratio, $\epsilon$, and leading-order term vary. Note how when $f(t)$ and $\kappa\Delta p/(2\pi)$ cancel each other, the asymmetric flow dominates. Moreover, the symmetry breaking becomes observable earlier for larger values of $\epsilon$, though we reiterate that our theory is formally valid only for $\epsilon\ll 1$.
We can get a first hint of the symmetry breaking mechanism observed in Figure 3 by considering when the asymmetric correction term $\epsilon w_{1}$ is of a similar size as the symmetric leading order solution $w_{0}$. Indeed, it is easy to verify that this will be the case when the pressure gradient approximately cancels out the forcing $f(t)$, such that the modulating coefficient $f(t)+\kappa\frac{\partial p_{0}}{\partial s}$ in (20a) is close to zero. This is visualized in Figure 4, where we plot the velocity $w=w_{0}+\epsilon w_{1}$ for several values of $\epsilon$ and $f(t)+\kappa\frac{\partial p_{0}}{\partial s}$, observing an asymmetric profile when $f(t)+\kappa\frac{\partial p_{0}}{\partial s}\ll 1$.
From the velocities (20), we may compute the flux, noting that the asymmetric components of $w_{1}(r,\theta,s,t)$ integrate to zero because of the $\cos\theta$ term:
$Q_{0}=-\frac{2\pi}{16}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)a(s)^{4},$ (21a)
@@
$I_{4}Q_{0}=-\frac{\pi}{8}\left(2\pi f(t)+\kappa\Delta p_{0}\right),$ (22a)
$I_{4}Q_{1}=-\frac{\pi}{8}\Delta p_{1},$ (22b)
where we have defined $I_{4}=\int_{0}^{2\pi}a(s)^{-4}~{}\mathrm{d}s$, and used the fact that $\Delta p_{1}=\Delta\bar{p}_{1}$.
@@
To connect the flow to the cupula displacement, we evaluate the flux using the velocity $w$ at the cupula’s location, where the fluid velocity must equal the velocity of the cupula: $w=\frac{\partial\eta}{\partial t}$. Therefore, we may write $Q_{0}=2\pi\int_{0}^{a_{0}}r\frac{\partial\eta_{0}}{\partial t}\mathrm{d}r$ where $a_{0}=a(0)$. Moreover, substituting the pressure jump using the plate equations (17d) and (18d) yields the following equations for the displacement $\eta$:
$I_{4}\int_{0}^{a(0)}r\frac{\partial\eta_{0}}{\partial t}\mathrm{dr}=-\frac{1}{16}\left[2\pi f(t)+\kappa\left(\nabla^{4}\eta_{0}-b^{-1}\nabla^{2}\eta_{0}+\varrho f(t)\right)\right],$ (23a)
$I_{4}\int_{0}^{a(0)}r\frac{\partial\eta_{1}}{\partial t}\mathrm{dr}=-\frac{1}{16}\kappa\left(\nabla^{4}\eta_{1}-b^{-1}\nabla^{2}\eta_{1}+\varrho r\cos\theta f(t)\right).$ (23b)
(23) are equations that completely determine $\eta_{0}$ and $\eta_{1}$ when the Stokes number is negligible. Furthermore, as $\varrho\ll 1$, we may neglect it in (23). We may solve (23) by writing the deflection as $\eta=\Delta p(t)\bar{\eta}(r)$, where $\bar{\eta}$ satisfies $\nabla^{4}\bar{\eta}-b^{-1}\nabla^{2}\bar{\eta}=1$, from which we obtain a pair of ordinary differential equations for $\Delta p(t)$:
@@
$\alpha(b)\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-\frac{1}{16I_{4}}\left(2\pi f(t)+\kappa\Delta p_{0}\right),$ (24a)
@@
$\alpha(b)\frac{\mathrm{d}\Delta p_{1}}{\mathrm{d}t}=-\frac{1}{16I_{4}}\kappa\Delta p_{1}$ (24b)
where we have defined $\alpha(b)=\int_{0}^{a_{0}}r\bar{\eta}(r)\mathrm{d}r$. If we consider the cupula’s edges to be clamped, the solution for $\bar{\eta}$ is given by:
$\bar{\eta}(r)=\frac{b}{4}(a_{0}^{2}-r^{2})+\frac{a_{0}b^{3/2}}{2I_{1}(b^{-1/2}a_{0})}\left[I_{0}\left(b^{-1/2}r\right)-I_{0}\left(b^{-1/2}a_{0}\right)\right],$ (25)
with $I_{k}(\cdot)$ the modified Bessel function of the first kind and $k$th order; we then also have
$\alpha(b)=\int_{0}^{a_{0}}r\bar{\eta}(r)\mathrm{d}r=\frac{a_{0}^{2}b^{2}}{2}+\frac{a_{0}^{4}b}{16}-\frac{a_{0}^{3}b^{3/2}I_{0}\left(a_{0}b^{-1/2}\right)}{4\,I_{1}\left(a_{0}b^{-1/2}\right)}.$ (26)
We remark that when no pretension exists in the cupula $\alpha(b\rightarrow\infty)$ is well-defined and finite, and in particular $\alpha\rightarrow a_{0}^{6}/384$, so that $\alpha$ remains well defined in the plate limit. (24b) has an exponential solution, but the condition of homogeneous initial conditions implies that $\Delta p_{1}(0)=0$, from which we conclude $\Delta p_{1}(t)\equiv 0$. Therefore, there is no first order correction to the cupular pressure jump when $\mathrm{St}\ll 1$ and $\varrho\ll 1$, and the leading order pressure pressure jump
$\Delta p_{0}=-\frac{\pi}{8\alpha(b)I_{4}}\int_{0}^{t}f(\tau)e^{-\frac{\kappa}{16I_{4}\alpha(b)}(t-\tau)}\mathrm{d}\tau,$ (27)
is correct to $\mathcal{O}(\epsilon^{2})$.
@@ 3.1.1 Expressions for the velocities
Once the pressure jump is known from (24a), the leading order pressure gradient may be computed from
$\frac{\partial p_{0}}{\partial s}=\frac{\Delta p_{0}}{I_{4}a(s)^{4}}+\frac{1}{\kappa}\left[\frac{2\pi}{I_{4}a(s)^{4}}-1\right].$ (28)
@@
$w_{0}$ $=\frac{\pi}{2I_{4}a(s)^{4}}\left[f(t)+\frac{\kappa\Delta p_{0}}{2\pi}\right][r^{2}-a(s)^{2}],$ (29a)
$w_{1}$ $=\frac{r(r^{2}-a(s)^{2})\cos\theta}{16I_{4}a(s)^{4}}\left[-3\kappa\Delta p_{0}+(-6\pi+4I_{4}a(s)^{4})f(t)\right].$ (29b)
@@
$u_{0}$ $=\frac{\pi}{2I_{4}a(s)^{5}}\left[f(t)+\frac{\kappa\Delta p_{0}}{2\pi}\right]r[r^{2}-a(s)^{2}]\frac{\mathrm{d}a}{\mathrm{d}s},$ (30a)
@@
$u_{1}$ $=\frac{2f(t)\left[2I_{4}a(s)^{4}+\pi\right]+\kappa\Delta p_{0}}{16I_{4}a(s)^{5}}r^{2}\left[r^{2}-a(s)^{2}\right]\cos\theta\frac{\mathrm{d}a}{\mathrm{d}s},$ (30c)
$v_{1}$ $=-\frac{5\left[2f(t)\left(2I_{4}a(s)^{4}+\pi\right)+\kappa\Delta p_{0}\right]}{16I_{4}a(s)^{5}}r^{2}\left[r^{2}-a(s)^{2}\right]\sin\theta\frac{\mathrm{d}a}{\mathrm{d}s}.$ (30d)
Note that for a perfect i.e. uniform torus, $a^{\prime}(s)=0$ and there are no radial or azimuthal velocities.
@@ 3.2 Deformation regimes
$\Delta p\sim-\frac{\pi}{8\alpha(b)I_{4}}\int_{0}^{t}f(\tau)\mathrm{d}\tau=-\frac{\pi}{8I_{4}\alpha(b)}F(t)\quad\text{as}\quad\kappa\to 0.$ (31)
@@
$\Delta p\sim-\frac{2\pi}{\kappa}f(t)\quad\text{as}\quad\kappa\to\infty.$ (32)
In physical terms, the two results in (31) and (32) represent qualitatively distinct regimes for the response of the cupular displacement to the imposed rotation of the canal: for a soft cupula ($\kappa\ll 1$), the pressure difference across the cupula, and thus the cupula deformation, is proportional to the angular velocity of the imposed rotation, $F(t)$, while for a stiff cupula ($\kappa\gg 1$), the deformation instead follows the angular acceleration $f(t)=F^{\prime}(t)$. Since the cupular deformation is thought to be what is detected by the nerve cells in the cupula, this suggests that the cupula can detect either the angular velocity or the angular acceleration to which it is subject — depending on the value of $\kappa$.
In the latter case, it is easy to verify that the leading order radial and axial velocities suffer a cancellation, as the prefactor of the leading order velocity (20a) is $f(t)+\kappa\frac{\Delta p_{0}}{2\pi}=\mathcal{O}\left(\frac{1}{\kappa}\right)$, and the asymmetric order $\epsilon$ correction $w_{1}(r,\theta,s,t)$ dominates for any $\epsilon$ provided $\kappa$ is sufficiently large. This cancellation accounts for the behaviour observed in Figure 3: as the Young’s modulus is increased, making the cupula stiffer to the point that $\kappa\gg 1$, the flow ceases to be symmetric.
@@
To estimate the critical of $\kappa$ where the transition occurs, we may write $\Delta p_{0}=Ae^{2\pi it}+c.c.$ and $f(t)=Be^{2\pi it}+c.c.$, with $A,B\in\mathbb{C}$,and seek the range of $\kappa$ for which $\Delta p$ is in phase with $f$. Direct substitution into (24a) yields
$A=-\frac{\pi}{8I_{4}}\frac{B}{2\pi\alpha(b)i+\kappa/(16I_{4})}.$ (33)
Therefore, the transition change occurs at $\kappa_{c}=32I_{4}\alpha(b)$. For $\kappa<\kappa_{c}$ the phase difference between $A$ and $B$ is more than $\frac{\pi}{4}$, and it will be less than $\frac{\pi}{4}$ for $\kappa>\kappa_{c}$. When the phase difference is small the response is roughly in phase with the forcing $f(t)$, and vice-versa.

### v2

@@ 3.1 Expanded solution
$w_{0}(r,s,t)=-\frac{1}{4}\left(\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}\right)(a(s)^{2}-r^{2}),$ (23a)
$w_{1}(r,\theta,s,t)=-\frac{1}{4}\frac{\partial\bar{p}_{1}}{\partial s}(a(s)^{2}-r^{2})-\frac{1}{16}\left(\dot{\Omega}(t)-3\frac{\partial p_{0}}{\partial s}\right)r(a(s)^{2}-r^{2})\cos\theta.$ (23b)
[…]
Figure 4: Velocity profiles predicted by (23) as the torus aspect ratio, $\epsilon$, and leading-order term vary. Note how when $\dot{\Omega}(t)$ and $\kappa\Delta p/(2\pi)$ cancel each other, the asymmetric flow dominates. Moreover, the symmetry breaking becomes observable earlier for larger values of $\epsilon$, though we reiterate that our theory is formally valid only for $\epsilon\ll 1$. Note that the horizontal axis may be interpreted as the phase difference between $\dot{\Omega}$ and $\Delta p$.
We can get a first hint of the symmetry breaking mechanism observed in Figure 3 by considering when the asymmetric correction term $\epsilon w_{1}$ is of a similar size as the symmetric leading order solution $w_{0}$. Indeed, it is easy to verify that this will be the case when the pressure gradient approximately cancels out the forcing $\dot{\Omega}(t)$, such that the modulating coefficient $\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}$ in (23a) is close to zero. This is visualized in Figure 4, where we plot the velocity $w=w_{0}+\epsilon w_{1}$ for several values of $\epsilon$ and $\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}$, observing an asymmetric profile when $\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}\ll 1$. We note that the above solutions are only valid far from the cupula, in the slender regions of the SCC ($s\gg\epsilon)$, and thus we refer to (23) as the outer solution. Closer to the cupula, a boundary layer develops due to the leading order velocity having to adjust from the outer solution $w_{0}(r,s,t)$ to the cupular deflection profile $\partial\eta_{0}/\partial t$ as imposed by (20) (see the curved streamlines close to the cupula in figure 3). The velocity adjustment causes the aforementioned pressure jump $\Delta p_{1}^{\text{BL}}(r,t)$, which we compute in Appendix F using a matched asymptotic expansion:
$\begin{split}&\Delta p_{1}^{\text{BL}}(r,t)=-\frac{\pi}{I_{4}}\left(\dot{\Omega}(t)+\frac{\Delta p_{0}(t)}{2\pi}\right)f_{3}(\beta)\mathrm{Re}\left\{\sum_{n=1}^{\infty}\bar{A}_{n}J_{0}(\mu_{n}r)\right\},\\ &f_{3}(\beta)=\frac{5(1-\nu_{s})}{4\beta^{2}(12-\nu_{s})+10(1-\nu_{s})},\end{split}$ (24)
where $\bar{A}_{n}\in\mathbb{C}$ are constants independent of time and cupular thickness and $\mu_{n}\in\mathbb{C}$ are the solutions in the first quadrant of $J_{1}(z)^{2}=J_{0}(z)J_{2}(z)$ (for details see Davis, 1990). From the velocities (23), we may compute the flux, noting that the asymmetric components of $w_{1}(r,\theta,s,t)$ integrate to zero because of the $\cos\theta$ term:
$Q_{0}=-\frac{2\pi}{16}\left(\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}\right)a(s)^{4},$ (25a)
@@
$I_{4}Q_{0}=-\frac{\pi}{8}\left(2\pi\dot{\Omega}(t)+\Delta p_{0}\right),$ (26a)
$I_{4}Q_{1}=-\frac{\pi}{8}\Delta p_{1}^{\text{outer}},$ (26b)
where we have defined $I_{4}=\int_{0}^{2\pi}a(s)^{-4}~\mathrm{d}s$, and used the fact that $\Delta p_{1}^{\text{outer}}=\Delta\bar{p}_{1}^{\text{outer}}$. We remark that when the thickness of the cupula is appreciable, the integrals should be performed from $s=\epsilon\beta/2$ to $s=2\pi-\epsilon\beta/2$, leading to a slight modification to (26), as discussed in Appendix E.
@@
To connect the flow to the cupula displacement, we evaluate the flux using the velocity $w$ at the cupula, where the fluid velocity must equal the velocity of the cupula: $w=\frac{\partial\eta}{\partial t}$. Therefore, we may write $Q_{i}=\int_{0}^{2\pi}\int_{0}^{a_{0}}r\frac{\partial\eta_{i}}{\partial t}\,\mathrm{d}r\mathrm{d}\theta$ where $a_{0}=a(0)$. In Appendix B.2 we show that the first order solid problem may be solved as
$\eta_{1}(r,\theta,t)$ $=\frac{\Delta p_{1}^{\text{outer}}(t)}{\kappa}\bar{\eta}_{0}(r)+\frac{\rho_{s}}{\rho\kappa}\dot{\Omega}(t)\bar{\eta}^{\text{euler}}_{1}(r)$ (27)
$+\frac{\rho_{s}\Omega_{0}\mathcal{T}}{\rho\kappa}\Omega(t)^{2}\bar{\eta}^{\text{centrif}}_{1}(r)\cos\theta-\frac{\pi}{I_{4}\kappa}\left(\dot{\Omega}(t)+\frac{\Delta p_{0}(t)}{2\pi}\right)f_{3}(\beta)\bar{\eta}^{\text{BL}}_{1}(r),$ (28)
where $\bar{\eta}_{0}$ was defined in (19) and other $\bar{\eta}_{1}$ are given in Appendix B.2. Computing the flux we find
$Q_{0}=\frac{1}{\kappa}\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}\int_{0}^{2\pi}\int_{0}^{a(0)}r\bar{\eta}_{0}(r;\beta)\,\mathrm{d}r\,\mathrm{d}\theta=\frac{2\pi\alpha_{0}(\beta)}{\kappa}\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t},$ (29a)
@@
$Q_{1}=\frac{2\pi}{\kappa}\left[\frac{\mathrm{d}\Delta p_{1}^{\text{outer}}}{\mathrm{d}t}\alpha_{0}(\beta)+\frac{\rho_{s}}{\rho}\ddot{\Omega}(t)\alpha_{1}^{\text{euler}}(\beta)-\frac{\pi}{I_{4}}\left(\ddot{\Omega}(t)+\frac{1}{2\pi}\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}\right)f_{3}(\beta)\alpha_{1}^{\text{BL}}(\beta)\right],$ (29b)
where the $\alpha$ factors are
$\alpha_{0}(\beta)\equiv\int_{0}^{a(0)}r\bar{\eta}_{0}(r;\beta)\mathrm{d}r=\frac{1-\nu_{s}^{2}}{32\beta^{3}}+\frac{(12-\nu_{s})(1+\nu_{s})}{80\beta},$ (30)
and similarly for the other contributions, with explicit forms given in §B.2. The centrifugal term integrates to zero, and we substitute the flux terms $Q_{0}$, $Q_{1}$ into (26) from which we obtain a pair of ordinary differential equations for $\Delta p(t)$:
@@
$\frac{\alpha_{0}(\beta)}{\kappa}\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-\frac{1}{16I_{4}}\left(2\pi\dot{\Omega}(t)+\Delta p_{0}\right),$ (31a)
$\frac{\alpha_{0}(\beta)}{\kappa}\frac{\mathrm{d}\Delta p_{1}^{\text{outer}}}{\mathrm{d}t}=-\frac{1}{16I_{4}}\Delta p_{1}^{\text{outer}}-\frac{\rho_{s}}{\rho}\frac{\alpha_{1}^{\text{euler}}(\beta)}{\kappa}\ddot{\Omega}(t)+\frac{\pi}{I_{4}\kappa}\alpha_{1}^{\text{BL}}(\beta)f_{3}(\beta)\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{\Omega}(t)+\frac{\Delta p_{0}}{2\pi}\right).$ (31b)
The boundary layer forcing term can be written in terms of $\Delta\ddot{p}_{0}$ using (31a), but it will be more useful to leave it in terms of $\Delta p_{0}$. We may solve (31a) for any forcing via the integral
$\Delta p_{0}=-\frac{\pi\kappa}{8\alpha_{0}(\beta)I_{4}}\int_{0}^{t}\dot{\Omega}(\tau)e^{-\frac{\kappa}{16I_{4}\alpha_{0}(\beta)}(t-\tau)}\mathrm{d}\tau.$ (32)
@@ 3.1.1 Expressions for the velocities
Once the pressure jump is known from (31a), the pressure gradient may be computed from
$\frac{\partial p_{0}}{\partial s}=\frac{\Delta p_{0}}{I_{4}a(s)^{4}}+\dot{\Omega}(t)\left[\frac{2\pi}{I_{4}a(s)^{4}}-1\right],\quad\frac{\partial\bar{p}_{1}}{\partial s}=\frac{\Delta p_{1}^{\text{outer}}}{I_{4}a(s)^{4}}.$ (33)
@@
$w_{0}=\frac{\pi}{2I_{4}a(s)^{4}}\left[\Omega(t)+\frac{\Delta p_{0}}{2\pi}\right][r^{2}-a(s)^{2}],$ (34a)
$w_{1}=\frac{r(r^{2}-a(s)^{2})\cos\theta}{16I_{4}a(s)^{4}}\left[-3\Delta p_{0}+(-6\pi+4I_{4}a(s)^{4})\Omega(t)\right]+\frac{\Delta p_{1}^{\text{outer}}}{4I_{4}a(s)^{4}}(r^{2}-a(s)^{2}).$ (34b)
@@
$u_{0}$ $=\frac{\pi}{2I_{4}a(s)^{5}}\left[\Omega(t)+\frac{\Delta p_{0}}{2\pi}\right]r[r^{2}-a(s)^{2}]\frac{\mathrm{d}a}{\mathrm{d}s},$ (35a)
@@
$u_{1}$ $=\frac{2\Omega(t)\left[2I_{4}a(s)^{4}+\pi\right]+\Delta p_{0}}{16I_{4}a(s)^{5}}r^{2}\left[r^{2}-a(s)^{2}\right]\cos\theta\frac{\mathrm{d}a}{\mathrm{d}s}+\frac{\Delta p_{1}^{\text{outer}}}{4I_{4}a(s)^{5}}r(r^{2}-a(s)^{2})\frac{\mathrm{d}a}{\mathrm{d}s},$ (35c)
$v_{1}$ $=-\frac{5\left[2\Omega(t)\left(2I_{4}a(s)^{4}+\pi\right)+\Delta p_{0}\right]}{16I_{4}a(s)^{5}}r^{2}\left[r^{2}-a(s)^{2}\right]\sin\theta\frac{\mathrm{d}a}{\mathrm{d}s}.$ (35d)
Note that for a perfect, i.e. uniform torus, $a^{\prime}(s)=0$ and there are no radial or azimuthal velocities.
@@ 3.2 Deformation regimes
$\Delta p_{0}\sim-\frac{\pi\kappa}{8\alpha_{0}(\beta)I_{4}}\int_{0}^{t}\dot{\Omega}(\tau)\mathrm{d}\tau=-\frac{\pi\kappa}{8I_{4}\alpha_{0}(\beta)}\Omega(t)\quad\text{as}\quad\kappa\to 0.$ (36)
@@
$\Delta p_{0}\sim-2\pi\dot{\Omega}(t)+\frac{32I_{4}\pi\alpha_{0}(\beta)}{\kappa}\ddot{\Omega}\quad\text{as}\quad\kappa\to\infty.$ (37)
In physical terms, the two results in (36) and (37) represent qualitatively distinct regimes for the response of the cupular displacement to the imposed rotation of the canal: for a soft cupula ($\kappa\ll 1$), the pressure difference across the cupula, and thus the cupula deformation, is proportional to the angular velocity of the imposed rotation, $\Omega(t)$, while for a stiff cupula ($\kappa\gg 1$), the deformation instead follows the angular acceleration $\dot{\Omega(t)}$. Since the cupular deformation is thought to be what is detected by the nerve cells in the cupula, this suggests that the cupula can detect either the angular velocity or the angular acceleration to which it is subject — depending on the value of $\kappa$.
In the latter case, it is easy to verify that the leading order radial and axial velocities suffer a cancellation, as the prefactor of the leading order velocity (23a) is $\dot{\Omega}(t)+\frac{\Delta p_{0}}{2\pi}=\mathcal{O}\left(\frac{1}{\kappa}\right)$, and the asymmetric order $\epsilon$ correction $w_{1}(r,\theta,s,t)$ dominates for any $\epsilon$ provided $\kappa$ is sufficiently large (the symmetric component of $w_{1}$ also scales as $\mathcal{O}(\kappa^{-1}))$. This cancellation accounts for the behaviour observed in Figure 3: as the Young’s modulus is increased, making the cupula stiffer to the point that $\kappa\gg 1$, the flow ceases to be symmetric. This is intuitive, as a completely rigid cupula does not allow a net flux and hence the axisymmetric leading order flow must vanish.
@@
To estimate the critical value of $\kappa$ where the transition occurs, we may write $\Delta p_{0}=Ae^{2\pi it}+c.c.$ and $\dot{\Omega}(t)=Be^{2\pi it}+c.c.$, with $A,B\in\mathbb{C}$,and seek the range of $\kappa$ for which $\Delta p$ is in phase with $\Omega$. Direct substitution into (31a) yields
$A=-\frac{2\pi B}{32I_{4}\pi\alpha(\beta)i/\kappa+1}.$ (38)
Therefore, the transition change occurs at $\kappa_{c}=32\pi I_{4}\alpha(b)$. For $\kappa<\kappa_{c}$ the phase difference between $A$ and $B$ is more than $\frac{\pi}{4}$, and it will be less than $\frac{\pi}{4}$ for $\kappa>\kappa_{c}$. When the phase difference is small the response is roughly in phase with the forcing $\Omega(t)$, and vice-versa.